WEB GEOMETRY: PAST, PRESENT AND FUTURE

Professor V.V. Goldberg New Jersey Institute Of Technology

The following topics will be discussed:

1. What is Web Geometry?

2. Developments in Web Geometry (1930–present).

3. Webs and Local Analytic Quasigroups.

4. Webs and Nomography.

5. Linearization Problem for Webs.

6. Webs and Algebraic Geometry: Grassmannization and Algebraization Problems for ebs.

7. Webs and Surfaces of Translation.

8. Connections of Web Geometry with Other Fields.

9. Rank Problems for Webs